# Multi-fingered Robotic Hand Grasping in Cluttered Environments through Hand-Object Contact Semantic Mapping

### v2

@@ IV-D Qualitative Study with SOTA methods
Qualitatively, we compare the grasp generation performance of our method with current contact information-guided grasp generation approaches, namely UniGrasp [1] and GenDexGrasp [2], as shown in Fig. 8. Additionally, we compare our method with the SOTA approach for grasp generation in cluttered environments, HGC-Net [8], to evaluate the performance in complex scenarios.
UniGrasp [1] predicts contact points one by one, while the CoSe-CVAE utilizes a generative model to predict all contact points simultaneously. As shown in Fig. 8 (a), the contact points generated by UniGrasp are relatively sparse, and the method exhibits limited diversity in its outcomes. As the number of contact points increases, UniGrasp’s ability to account for the relationships between contact points diminishes, resulting in contact point predictions that fail to generate feasible grasp postures. CoSe-CVAE is more effective at considering the relationships between contact points, enabling it to generate diverse contact semantic maps.
@@
GenDexGrasp [2] estimates contact distance map to guide optimization-based grasp generation, but it is prone to sub-optimal solutions in global optimization. As shown in Fig. 8 (b), the generated grasp samples do not match the predicted contact distance map. In contrast, the contact semantic map in our methods provides more accurate guidance for grasp generation, as shown in Fig. 8 (c).
The planning results for cluttered environments with unknown objects, using both our approach and the HGC-Net [8], are presented in Fig. 8. In real-world cluttered scenarios, HGC-Net demonstrates instability when handling objects under collision conditions, often resulting in failed grasps. As illustrated in the Fig. 8 (c) and (d), the optimal grasps generated by our model consistently outperform those of HGC-Net, delivering higher-quality results.
Incorporating contact semantic information into the grasping process significantly avoids local optimal candidate with opposite grasp pose, enhances the accuracy of the outcomes, leading to more effective grasp generation. Our grasp evaluation model is capable of assessing grasps in cluttered environments, consistently identifying the optimal grasp.
TABLE II: Outcomes of real-world grasping experiments
Method Success Rate*(%)
household object YCB object Cluttered
GenDexGrasp [2] 62.7 58.7 -
UniGrasp [1] 71.3 70.7 -
HGC [8] 78.7 74.0 70.0
Ours 85.3 76.7 75.3
    Success rate is determined by executing 150 grasping attempts.
IV-E Quantitative Grasping Experiments
@@
We conduct experiments in single-object grasping and grasping from cluttered environments, employing both our proposed method and baselines. Using our method, the grasping average success rate in single-object scenes reaches 81.0%, surpassing other baseline approaches [2, 1, 8]. Our grasp detection methods enhance the accuracy of grasping candidate selection. Additionally, our evaluation model substantially aids in filtering out unfeasible grasps for cluttered scenes. In grasping experiments from cluttered scenes, our method achieves a 75.3% success rate.
Most grasping failures are due to object slippage and ineffective force closure. Errors in joint angles during grasp generation cause slippage, while inaccuracies during the process prevent proper force closure. Failures often occur on object handles, where force-closure areas are limited.
IV-F Grasping Evaluation for Different Multi-fingered Robotic Hands
To verify the generalization capability of our grasp evaluation model across different robotic hands, we performed real-world grasping experiments in cluttered scenes using both known hand, DLR-HIT II hand, and unknown robotic hand, four-fingered LEAP hand [47]. The results of the grasp generation with both DLR-HIT II hand and LEAP hand using same contact semantic maps are shown in the Fig 9. The average success rate of unknown hand achieves 74.0% in 50 grasping attempts.
Figure 9: Contact semantic map-guided grasp generation results using known DLR-HIT II hand and unknown LEAP hand.
V Conclusions and Future Work
@@
We propose a novel method for multi-fingered grasp generation in cluttered environments, leveraging semantic contact information. First, the CoSe-CVAE model predicts diverse contact semantic maps between the hand and the object from the object’s point cloud. The grasp detection method then estimates grasp poses based on these semantic contact maps. Furthermore, our proposed grasp evaluation network utilizes both scene and robotic hand point clouds to predict grasp quality, selecting the optimal grasp in cluttered scenes. In this work, we also propose a pipeline for generating multi-modal multi-fingered robotic hand grasping dataset.
@@
Qualitative real-world experiments demonstrate that our CoSe-CVAE model can reliably generate hand-object contact semantic information, significantly enhancing the stability of grasp generation based on contact information, outperforming SOTA methods [2, 1]. By incorporating the geometric characteristics of the robotic hand, the proposed grasp evaluation model can more effectively assess the grasp quality of multi-fingered hands in cluttered environments, outperforming [8]. Quantitative comparisons in real-world experiments further show that our method achieves a higher grasp success rate than these SOTA methods, with an average success rate of 81.0% in single-object scenarios and 75.3% in multi-object scenarios. Additionally, using an unknown robotic hand, we validate that our method achieves an average success rate of 74.0%.
In future work, we will explore how to leverage contact information to assist in complex and dexterous manipulations with robotic hands.

### v3

@@ IV-D Contact Information-Guided Grasping Generation
Qualitatively, we compare the grasp generation performance of our method with current contact information-guided grasp generation methods, namely GenDexGrasp [2], ContactGen [3], and UniGrasp [1], as shown in Fig. 8.
@@
GenDexGrasp [2] estimates contact distance map to guide optimization-based grasp generation, but it is prone to sub-optimal solutions in global optimization and suffers from semantic ambiguity. Specifically, the generated grasps do not always align with the predicted contact semantic information, leading to inconsistencies between the grasp poses and the intended contact regions. As shown in Fig. 8 (a), the generated grasp samples do not match the predicted contact distance map. In contrast, the CoSe-CVAE provides more accurate guidance for grasp generation.
ContactGen [3] generates a sequence of contact maps for human grasp synthesis, including a contact map, a part map, and a direction map. However, the complexity of this model leads to cumulative errors across all predicted maps, which propagate to the final grasp result. Moreover, discrepancies in size and joint design between human and robotic hands often prevent multi-fingered from replicating human grasp poses accurately. As a result, ContactGen struggles to generalize well to robotic hands, as illustrated in Fig. 8 (b). In contrast, our approach utilizes a single contact semantic map, effectively improving stability in multi-finger robotic grasp generation.
UniGrasp [1] predicts contact points one by one, while the CoSe-CVAE utilizes a generative model to predict all contact points simultaneously. As shown in Fig. 8 (c), the contact points generated by UniGrasp are relatively sparse, and the method exhibits limited diversity in its outcomes. As the number of estimated contact points increases, UniGrasp’s ability to account for the relationships between contact points diminishes, resulting in contact point predictions that fail to generate feasible grasp postures. Moreover, representing with sparse contact points introduces additional complexity, as a single grasp sample corresponds to multiple configurations with contact points, increasing the difficulty of model training. CoSe-CVAE is more effective at considering the relationships between contact points, enabling it to generate diverse contact semantic maps.
Incorporating contact semantic map into the grasping process significantly improves stability of grasp generation, enhances semantic consistency.
IV-E Unknown Multi-fingered Robotic Hand Grasp Generation
To verify the generalization capability of SOTA methods [2, 3, 1] and our CoSe-CVAE across different robotic hands, we estimate grasp candidates using known hand, DLR-HIT II hand, and unknown robotic hand, four-fingered LEAP hand [53] from identical contact maps. The results are shown in Fig. 8 (d) and (e). Compared to SOTA methods, CoSe-CVAE provides more stable and precise guidance for grasp pose estimation of unknown robotic hands.
IV-F Grasping from Cluttered Scenes
To evaluate the performance in cluttered scenarios, we perform comparison experiments using our method and the SOTA approach, HGC-Net [11]. The planning results for cluttered environments with unknown objects are presented in Fig. 9. The optimal grasps generated by our model consistently outperform those of HGC-Net, delivering higher-quality results. In cluttered real-world scenarios, HGC-Net often predicts grasp poses that result in unintended collisions with objects. This premature finger contact can lead to grasp failures, reducing overall grasp success rates. By incorporating a contact semantic map between the perception and grasping processes, our pipeline is capable of assessing grasps in cluttered environments, consistently identifying the optimal grasp.
Figure 9: Inference results for cluttered scenes. (a) Grasp candidates based on HGC-Net [11]. (b) Positive grasp candidates based on our methods.
IV-G Quantitative Grasping Experiments
We conduct comparison experiments in grasping a single object using known and unknown hands, as well as grasping from cluttered environments using a known hand. For each setting, we perform $150$ grasp attempts per method. A grasp is considered successful if the robotic hand securely lifts the object and maintains a stable grip for at least two seconds. Any grasp predicted as successful, but resulting in a collision, will be automatically considered a failure. The success rate is determined by calculating the proportion of successful grasps out of the total number of attempts. The quantitative results are shown in Tab. II.
TABLE II: Quantitative results of real-world grasping experiments
Method Success Rate(%)
Single-Object Scene Single-Object Scene Cluttered Scene
Known Hand Unknown Hand Known Hand
Household YCB Household YCB Household YCB
GenDexGrasp [2] 62.7 58.7 60.7 55.3 - -
ContactGen [3] 64.7 63.3 62.0 59.3 - -
UniGrasp [1] 71.3 70.7 70.7 64.7 - -
HGC [11] 78.7 74.0 - - 70.7 70.0
[…]
IV-G1 Single-Object Grasping using Known Hand
@@
Using our method, the average grasping success rate reaches 81.0% in single-object scenes, surpassing other baseline approaches [2, 1, 11, 3]. Our grasp detection methods improve the accuracy of grasping candidate selection.
IV-G2 Contact Information-Guided Grasping using Unknown Hand
We conduct comparison experiments of real-world robotic grasping using the unknown hand. The average success rate of our pipeline reaches 76.7%, outperforming the SOTA methods [2, 3, 1]. We also conduct real-world grasping experiments from cluttered experiments using the unknown hand based on our approach. The average success rate achieves 74.0%.
IV-G3 Grasping from Cluttered Scenes
In grasping experiments from cluttered scenes, our method achieves a 76.7% average success rate. Our evaluation model substantially aids in filtering out unfeasible grasps for cluttered scenes.
Figure 10: Examples of successful and failed grasping trials.
IV-H Grasping Failure Analysis
Most grasping failures are due to collisions with the surrounding environment, object slippage and ineffective force closure, as shown in Fig. 10. Our method results in fewer grasping failures due to collisions with the surrounding environment compared to HGC-Net [11]. In grasping experiments from cluttered scenes, our collision failure rate is 4.0%, whereas HGC-Net’s is 10.7%, representing a 62.5% reduction in collision failures. Errors in joint angles during grasp generation cause slippage, while inaccuracies during the process prevent proper force closure. This issue can be mitigated through tactile-based manipulation, which we will explore in future work.
V Limitations and Future Work
The current dataset has a much higher number of pinch grasps compared to other grasp types, as the data generation algorithm [24] cannot explicitly generate candidates with specific grasp types. Future work will incorporate contact information for grasp type-aware generation. Secondly, our focus is on grasp pose generation, with robot trajectories in experiments determined via path planning. We leave policy-based trajectory generation as a direction for future work.
VI Conclusions
@@
We propose a novel semantic contact information-guided grasp generation method for multi-fingered robotic hands in single-object and cluttered environments. First, the CoSe-CVAE model predicts diverse contact semantic maps between the hand and the object from the object’s point cloud. The grasp detection method then estimates grasp poses based on these semantic contact maps. Furthermore, our proposed grasp evaluation network PointNetGPD++ utilizes both scene and robotic hand point clouds to predict grasp quality, selecting the optimal grasp in cluttered scenes.
@@
Qualitative real-world experiments demonstrate that our CoSe-CVAE model can reliably generate hand-object contact semantic information, significantly enhancing the stability of grasp generation based on contact information using known and unknown hands, outperforming SOTA methods [2, 1, 3]. By incorporating the geometric characteristics of the robotic hand, the proposed grasp evaluation model can more effectively assess the grasp quality of multi-fingered hands in single-object and cluttered environments, outperforming SOTA methods [11, 3, 2, 1]. Quantitative comparisons in real-world experiments further show that our method achieves a higher grasp success rate than these SOTA methods, with an average success rate of 81.0% in single-object scenarios and 76.7% in multi-object scenarios. Additionally, average success rate of grasp experiments using unknown robotic hand reaches 76.7% in single-object scenes, surpassing SOTA methods [2, 1, 3] by at least 9.0%.